A 25.00 mL sample of nitric acid was titrated with 0.6440 M strontium hydroxide. A total of 10.22 mL of the base was required to reach the equivalence point.


a) Write a balanced molecular equation for this reaction?


b) What was the concentration of the nitric acid?


1
Expert's answer
2021-05-25T07:51:48-0400

"a)\ 2HNO_3 + Sr(OH)_2 \to Sr[(NO)_3]_2 + 2H_2O"


"b)\"Using, "\\dfrac{\u200bC_A\u200bV_A\u200b\u200b}{C_B\u200bV_B} =\\dfrac{n_A}{n_B }"


"\dfrac{C_A ×25}{0.6440×10.22}=\dfrac21"


"3.79842 C_A= 2"


"C_A = 0.53M"